Benoit Company produces three products—A, B, and C. Data concerning the three products follow (per unit):

Product

A B C
Selling price $

72.00

$ 60.00 $ 62.00
Variable expenses:
Direct materials 21.60 18.00 9.00
Other variable expenses 21.60 27.00 34.40
Total variable expenses 43.20 45.00 43.40
Contribution margin $ 28.80 $ 15.00 $ 18.60
Contribution margin ratio 40 % 25 % 30 %

The company estimates that it can sell 1,000 units of each product per month. The same raw material is used in each product. The material costs $3 per pound with a maximum of 6,600 pounds available each month.

Required:

1. Calculate the contribution margin per pound of the constraining resource for each product.

2. Which orders would you advise the company to accept first, those for A, B, or C? Which orders second? Third?

3. What is the maximum contribution margin that the company can earn per month if it makes optimal use of its 6,600 pounds of materials?
Answer #1
1
A B C
Contribution margin 28.80 15.00 18.60
Divide by Direct materials per unit 7.2 6 3
Contribution margin per pound of the
constraining resource
4 2.5 6.2
2
First - Product C
Second - Product A
Third - Product B
3
Units of Product C = 1000
Remaining raw materials = 6600-(1000*3)= 3600
Units of Product A = 3600/7.2= 500
Maximum contribution margin 33000 =(1000*18.6)+(500*28.8)